# Neural timescales from a computational perspective

### v1

@@ 5.4 Overlooked timescales and outlook
Much of systems and cognitive neuroscience has been based on a specific set of tasks, often with a trial structure between 1 to 10 seconds long, which artificially limits observed neural and behavioral timescale to a small range of the spectrum. For instance, spontaneous animal behavior in natural environments exhibits correlations over a long range of timescales that cannot be measured in common experimental settings [171, 172, 173]. On the other hand, processes such as the release of synaptic vesicles occur on a sub-second timescale [174] that are not detectable with many measurement tools. Furthermore, different physiological processes may be involved in shaping neural timescales that are often ignored when studying the mechanisms underlying neural timescales. For example, astrocytes have slow intracellular calcium fluctuations that may mediate slow neural modulations [175]. Astrocytes may also play a significant role in neural computations, as recently was shown that neuron–astrocyte networks can perform similar computations to complex machine learning models like Transformers [176]. How processes across different timescales interact may be of tremendous interest: does sub-second dynamics of synaptic vesicles play a role in shaping how, for example, memory engrams are formed over days and months, beyond the mean activity? On the other hand, do longer timescale fluctuations beyond a designated static state like age or disease conditions, such as circadian or seasonal rhythms, impact moment-to-moment neural dynamics? To investigate these questions, we likely need to leverage a combination of different experimental paradigms in the lab and naturalistic settings [171, 177, 178], together with long-term and high temporal resolution neural recording [179, 9], as well as novel data analysis and modeling tools to parse the interaction between and across the different orders of magnitude timescales.
@@
The authors thank Ana Manea and Jan Zimmermann for providing the timescales of fMRI data (Fig. 2a), Bradley Voytek and the speakers and participants of the Cosyne 2022 workshop on “Mechanisms, functions, and methods for diversity of neuronal and network timescales” for valuable discussions, and Scidraw, Wikipedia and Pixel perfect from Flaticon for providing the basis icons for some of the illustrations.

### v2

@@ 5.4 Overlooked timescales and outlook
Much of systems and cognitive neuroscience has been based on a specific set of tasks, often with a trial structure between 1 to 10 seconds long, which artificially limits observed neural and behavioral timescale to a small range of the spectrum. For instance, spontaneous animal behavior in natural environments exhibits correlations over a long range of timescales that cannot be measured in common experimental settings [210, 211, 212]. On the other hand, processes such as the release of synaptic vesicles occur on a sub-second timescale [213] that are not detectable with many measurement tools. Furthermore, different physiological processes may be involved in shaping neural timescales that are often ignored when studying the mechanisms underlying neural timescales. For example, astrocytes have slow intracellular calcium fluctuations that may mediate slow neural modulations [214]. Astrocytes may also play a significant role in neural computations, as was recently shown that neuron–astrocyte networks can perform similar computations to complex machine learning models like Transformers [215]. How processes across different timescales interact may be of tremendous interest: does sub-second dynamics of synaptic vesicles play a role in shaping how, for example, memory engrams are formed over days and months, beyond the mean activity? On the other hand, do longer timescale fluctuations beyond a designated static state like age or disease conditions, such as circadian or seasonal rhythms, impact moment-to-moment neural dynamics? To investigate these questions, we likely need to leverage a combination of different experimental paradigms in the lab and naturalistic settings [210, 216, 217], together with long-term and high temporal resolution neural recording [218, 41], as well as novel data analysis and modeling tools to parse the interaction between timescales across different orders of magnitude.
@@
We thank Ana Manea and Jan Zimmermann for providing the timescales of fMRI data (Fig. 2a), Bradley Voytek and the speakers and participants of the Cosyne 2022 workshop on “Mechanisms, functions, and methods for diversity of neuronal and network timescales” for valuable discussions, and Scidraw (Federico Claudi and Andrew Hardaway), Wikipedia and Pixel perfect from Flaticon for providing the basis icons for some of the illustrations, and International Max Planck Research School for the Mechanisms of Mental Function and Dysfunction (IMPRS-MMFD) and Joachim Herz Foundation for their support.
Author contributions
RZ and RG conceptualized the overall structure of the review, wrote the first draft, and prepared the figures. All authors discussed, revised, and wrote the final version of the paper.